# Federated Multi-armed Bandits Under Byzantine Attacks

### v2

@@ Abstract
Multi-armed bandits (MAB) is a sequential decision-making model in which the learner controls the trade-off between exploration and exploitation to maximize its cumulative reward. Federated multi-armed bandits (FMAB) is an emerging framework where a cohort of learners with heterogeneous local models play an MAB game and communicate their aggregated feedback to a server to learn a globally optimal arm. Two key hurdles in FMAB are communication-efficient learning and resilience to adversarial attacks. To address these issues, we study the FMAB problem in the presence of Byzantine clients who can send false model updates threatening the learning process. We analyze the sample complexity and the regret of $\beta$-optimal arm identification. We borrow tools from robust statistics and propose a median-of-means (MoM)-based online algorithm, Fed-MoM-UCB, to cope with Byzantine clients. In particular, we show that if the Byzantine clients constitute less than half of the cohort, the cumulative regret with respect to $\beta$-optimal arms is bounded over time with high probability, showcasing both communication efficiency and Byzantine resilience. We analyze the interplay between the algorithm parameters, a discernibility margin, regret, communication cost, and the arms’ suboptimality gaps. We demonstrate Fed-MoM-UCB’s effectiveness against the baselines in the presence of Byzantine attacks via experiments.
Index Terms:
@@
I Introduction
Multi-armed bandits (MAB) model sequential decision-making under uncertainty where the goal is to maximize the collected reward under noisy feedback over a sequence of rounds from $1$ up to $T$ [1]. In its simplest form, the learner interacts with $K$ arms with unknown reward distributions. At each round, the learner pulls an arm and collects its immediate reward. In order to maximize the collected reward, the learner should intricately balance information collection and utilization. This led to the development of a series of algorithms that encoded exploration and exploitation either explicitly or implicitly in their arm selection mechanisms. Examples of the former include explore-then-commit [2] and interleaved exploration-exploitation [3] style algorithms. Examples of the latter include Thompson sampling [4] and upper confidence bound [5] strategies. Instead of measuring how high the rewards of these algorithms are, researchers have adopted a more interpretable benchmark, namely, how well these algorithms perform with respect to the best possible sequence of arms. This performance metric is known as the regret. It is known that any reasonable algorithm that performs well under a wide range of reward distributions should suffer at least $\Omega(\log T)$ expected regret [6]. Remarkably, many variants of the aforementioned algorithms succeed in achieving $O(\log T)$ expected regret, confirming their time order optimality.
@@ I Introduction
In the literature, MAB models have been particularly instrumental in transforming recommender systems, where they address the challenge of providing personalized content to users by continuously learning from user interactions. A notable example is [10], where a contextual bandit approach is employed to enhance news article recommendations. Other works highlighting the success of MAB algorithms in recommendation systems include [11, 12]. Similarly, by skillfully matching tasks with the best-suited participants, MAB algorithms emerge as a powerful tool in crowd-sourcing [13, 14, 15, 16, 17, 18]. Some of them smartly assign tasks outside of participants’ usual expertise, uncovering hidden abilities and adjusting to skill changes. This method improves how tasks are assigned, making the process more efficient and successful.
@@
  • We formalize the FMAB framework with Byzantine clients (Section III).
  • We propose a median-of-means (MoM) based algorithm, Fed-MoM-UCB, and a trimmed mean based algorithm, Fed-Trim-UCB, to maintain robustness against the Byzantine clients (Section IV).
  • We show that utilizing a MoM estimator ensures the elimination of arms with suboptimality gaps smaller than a discernibility margin (Section V).
  • We derive $(\beta,\delta)$-PAC sample complexity results and a high probability $O(1)$ upper bound on the cumulative regret for our algorithm (Section V).
  • We provide guidelines and insights towards the parameter selection for the MoM estimator considering the trade-offs between the number of required clients, the cumulative regret, and the discernibility margin (Section V).
  • We conduct detailed experiments to compare Fed-MoM-UCB’s performance against Fed2-UCB in [21] and Fed-Trim-UCB under Byzantine attacks and to investigate the effects of different parameters on the performance (Section VI).
II Related Work
TABLE I: Comparison of different works on FMAB problems
@@ II Related Work
Byzantine-UCB [27] $\times$ $\times$ $\widetilde{O}(T^{3/4})$$\times$
@@
II-A Federated Learning
In recent years, FL has experienced significant advancements, with particular focus on addressing key challenges such as privacy and communication costs, as thoroughly explored in [29]. The work by [30] introduces MiTFed, an innovative framework that leverages blockchain technology and smart aggregation to enhance robustness against Distributed Denial of Service (DDoS) attacks, ensuring privacy and facilitating collaboration in a decentralized context. Another study by [31] tackles the issues of communication overhead and data heterogeneity—two major obstacles in FL as identified by [32]. The authors propose the COTAF algorithm, which preserves the convergence properties of local Stochastic Gradient Descent (SGD) while accommodating heterogeneous data across users. Additionally, [33] addresses data heterogeneity by quantifying the skewness of client data in a privacy-preserving manner, subsequently selecting clients with less skewed data using a bandit approach. Scalability of FL systems in edge networks is studied in [34], where a fully decentralized FL system is shown to achieve significant gains in scalability and adaptivity.
@@ II-A Federated Learning
II-B Federated Multi-armed Bandits
@@
FMAB problems share common challenges with FL, and several studies have sought to address various FL aspects within the FMAB framework. Table I provides a comparison of different FMAB studies, outlining the specific challenges each addresses. Notably, [21] was among the first to introduce the FMAB framework, with a focus on data heterogeneity and communication costs. Another significant work is by [27], which explores a federated linear bandit scenario involving Byzantine clients, albeit without considering data heterogeneity. Their main objective is to enhance the performance of each uncorrupted client by minimizing a metric they refer to as robust regret, which aggregates the regrets of all uncorrupted clients but does not account for communication costs. Their approach offers a fully personalized solution, positioning it at the other end of the spectrum compared to [21]. Similarly, [28] operates in a comparable environment where data heterogeneity manifests through the availability of arms to different clients, rather than varying arm means across clients. This study zeroes in on the best arm identification problem, securing high-probability identification outcomes. Recently, global reward maximization by utilizing partial distributed feedback is studied by [45] within the context of linear bandits, while also aiming differential privacy. Each study mentioned adopts a distinct approach, combining various aspects. In our work, we address both $(\beta,\delta)$-PAC BAI and regret minimization amidst heterogeneous data and Byzantine attacks, while also factoring in communication costs.
III Problem Formulation
III-A Federated MAB Setup
@@
We consider a federated learning system with a server and a pool of clients that can be recruited by the server. There are $K$ arms indexed by the set $[K]\coloneqq\{1,2,...,K\}$. Arm $k$ has an unknown global mean reward $\mu_{k}$. We call $\boldsymbol{\mu}:=(\mu_{k})_{k\in[K]}$ the global model. We denote the global optimal arm by $k^{*}\coloneqq\operatorname*{argmax}_{k\in[K]}\mu_{k}$ and assume that it is unique. Clients cannot directly interact with the global model, but instead, they interact with their local models, which are stochastic realizations of the global model. In the local model of client $m$, the local mean reward of arm $k$ is generated as $\mu_{m,k}=\mu_{k}+\xi_{m,k}$, where $\xi_{m,k}$ is the $\sigma_{c}$-subgaussian model noise. For an illustrative example of the model noise, consider a recommender system in which the server’s objective is to pinpoint the item enjoying the highest global popularity, aiming to maximize profits. Due to the server’s lack of direct access to global popularities, it relies on feedback from selected clients. Nonetheless, the popularities observed by clients at a local level may diverge from the global trends. The random variable $\xi_{m,k}$ represents this discrepancy. We assume that $\sigma_{c}$ is known by the server.
@@
$X_{m,k}(t)=\mu_{m,k}+\zeta_{t}=\mu_{k}+\xi_{m,k}+\zeta_{m,t}~{}.$ (1)
In the above display, $\zeta_{m,t}$ is the sampling noise that is drawn from a fixed $\sigma$-subgaussian distribution independent of other clients and rounds. We assume that $\sigma$ is known by the server. Given $\beta>0$ and $\delta\in(0,1)$, the server and the clients’ joint goal is to identify an arm whose global mean is within $\beta$ of the global optimal arm with at least $1-\delta$ probability. In the bandit literature, this is also known as $(\beta,\delta)$-PAC BAI [7].
Note that while $\zeta_{m,t}$ is sampled at each round $t$, $\xi_{m,k}$ is sampled only once at the beginning, and $\mu_{m,k}$ is fixed from thereon. Equivalently, we can view $X_{m,k}(t)$ as a sample from an unknown distribution $\nu_{m,k}$. We call $\boldsymbol{\nu}_{m}:=(\nu_{m,k})_{k\in[K]}$ the local model of client $m$. For different clients $m$ and $n$, we have $\mu_{m,k}\neq\mu_{n,k}$ in general. Since the local mean rewards $\mu_{m,k}$ are different from the global mean rewards $\mu_{k}$, a client $m$ may not agree with the global optimal arm, i.e., $k^{*}\notin\operatorname*{argmax}_{k\in[K]}\mu_{m,k}$. In such a case, it is a hopeless task to identify the global optimal arm by relying on the local observations of a single client. However, the structure of the model noise suggest that by recruiting many clients, and averaging over their samples, one should be able to get an accurate estimate of the global mean rewards. This idea is exploited in [21], where the interaction between the server and the clients proceeds in phases indexed by $p\geq 1$. At the end of each phase, clients send updates to the server, denoted by $U_{m,k}(p)$, which are estimates of the local mean rewards of arms. The server, then aggregates these updates to discard suboptimal arms and informs the clients. After this, the clients refine their updates by sampling the remaining arms in the next phase. This interaction continues until $k^{*}$ is identified.
III-B Byzantine Clients
@@
In our study, we focus on the challenge of identifying $\beta$-optimal arms, namely, those arms residing within a $\beta$-proximity of the global optimal arm. To address the robustness concern, we introduce a median-of-means (MoM) aggregator for the server, detailed further in Section IV. It will be impossible to learn $\beta$-optimal arms when all recruited clients are Byzantine. In order to obtain meaningful results, we make the following assumption on the ratio of the Byzantine clients. This assumption is standard in the robust statistics literature, where it is assumed that an upper bound on the fraction of corrupted samples is known [46, 47, 48, 28].
@@
III-C Regret
In addition to $\beta$-optimal arm identification, we also want our algorithm to avoid sampling $\beta$-suboptimal arms unnecessarily, until all $\beta$-suboptimal arms are eliminated. We measure the performance of the interaction protocol between the server and the clients by introducing the notion of robust federated regret. Let $\Delta_{k}:=\mu_{k^{*}}-\mu_{k}$ denote the suboptimality gap for arm $k$. For $\beta\geq 0$, let $\mathcal{K}_{\beta}:=\{k\in[K]~{}|~{}\Delta_{k}\leq\beta\}$ denote the set of $\beta$-optimal arms. Let $\neg\mathcal{K}_{\beta}\coloneqq[K]\setminus\mathcal{K}_{\beta}$ denote the set of $\beta$-suboptimal arms. We define our objective as finding an arm $k\in\mathcal{K}_{\beta}$. Note that when $\beta<\min_{k\in[K]\setminus\{k^{*}\}}\Delta_{k}$, this coincides with finding the optimal arm. Let $I_{t}^{m}$ denote the arm selected by client $m$ at round $t$. We denote by $\mathcal{M}_{t}^{\beta}\subseteq\mathcal{M}$ the set of clients who plays an arm $k\in\neg\mathcal{K}_{\beta}$ at round $t$. Robust federated regret defined below accounts for the cumulative loss of all clients with respect to $\beta$-optimal arms by considering the communication cost between the server and clients.

### v3

@@ Abstract
Multi-armed bandits (MAB) is a sequential decision-making model in which the learner controls the trade-off between exploration and exploitation to maximize its cumulative reward. Federated multi-armed bandits (FMAB) is an emerging framework where a cohort of learners with heterogeneous local models play a MAB game and communicate their aggregated feedback to a server to learn a globally optimal arm. Two key hurdles in FMAB are communication-efficient learning and resilience to adversarial attacks. To address these issues, we study the FMAB problem in the presence of Byzantine clients who can send false model updates threatening the learning process. We analyze the sample complexity and the regret of $\beta$-optimal arm identification. We borrow tools from robust statistics and propose a median-of-means (MoM)-based online algorithm, Fed-MoM-UCB, to cope with Byzantine clients. In particular, we show that if the Byzantine clients constitute less than half of the cohort, the cumulative regret with respect to $\beta$-optimal arms is bounded over time with high probability, showcasing both communication efficiency and Byzantine resilience. We analyze the interplay between the algorithm parameters, a discernibility margin, regret, communication cost, and the arms’ suboptimality gaps. We demonstrate Fed-MoM-UCB’s effectiveness against the baselines in the presence of Byzantine attacks via experiments.
{IEEEImpStatement}
Federated learning systems became ubiquitous in many real world applications, such as healthcare, recommender systems, and communications. A critical challenge in ensuring such systems’ robustness and communication efficiency involves dealing with malicious agents that send falsified updates. Our work develops a principled approach to Byzantine robust federated sequential decision-making by proposing a new federated multi-armed bandit problem. It not only characterizes the fundamental difficulty of finding the optimal arms in the presence of Byzantine clients but also rigorously bounds the sample complexity and regret of federated sequential decision-making. Thanks to its strong theoretical guarantees, our algorithm significantly improves the trustworthiness of federated sequential decision-making systems, paving the way for secure and resilient federated learning applications. Our fundamental theoretical results on learnability and Byzantine resilience, together with comprehensive experimental evaluations of the proposed algorithms, pave the way for the widespread adoption of secure and robust federated learning systems in the real world.
{IEEEkeywords}
@@
1 Introduction
[…]
Multi-armed bandits (MAB) model sequential decision-making under uncertainty where the goal is to maximize the collected reward under noisy feedback over a sequence of rounds from $1$ up to $T$ [1, 2, 3]. In its simplest form, the learner interacts with $K$ arms with unknown reward distributions. At each round, the learner pulls an arm and collects its immediate reward. In order to maximize the collected reward, the learner should intricately balance information collection and utilization. This led to the development of a series of algorithms that encoded exploration and exploitation either explicitly or implicitly in their arm selection mechanisms. Examples of the former include explore-then-commit [4] and interleaved exploration-exploitation [5] style algorithms. Examples of the latter include Thompson sampling [6] and upper confidence bound [7] strategies. Instead of measuring how high the rewards of these algorithms are, researchers have adopted a more interpretable benchmark, namely, how well these algorithms perform with respect to the best possible sequence of arms. This performance metric is known as the regret. It is known that any reasonable algorithm that performs well under a wide range of reward distributions should suffer at least $\Omega(\log T)$ expected regret [8]. Remarkably, many variants of the aforementioned algorithms succeed in achieving $O(\log T)$ expected regret, confirming their time order optimality.
@@ 1 Introduction
In the literature, MAB models have been particularly instrumental in transforming recommender systems, where they address the challenge of providing personalized content to users by continuously learning from user interactions. A notable example is [12], where a contextual bandit approach is employed to enhance news article recommendations. Other works highlighting the success of MAB algorithms in recommendation systems include [13, 14]. Similarly, by skillfully matching tasks with the best-suited participants, MAB algorithms emerge as a powerful tool in crowd-sourcing [15, 16, 17, 18, 19, 20]. Some of them smartly assign tasks outside of participants’ usual expertise, uncovering hidden abilities and adjusting to skill changes. This method improves how tasks are assigned, making the process more efficient and successful. Recently, bandit algorithms have also been used for joint device activity detection and data decoding in random access with massive Internet-of-Things devices [21].
@@
  • We propose a median-of-means (MoM) based algorithm, Fed-MoM-UCB, and a trimmed mean based algorithm, Fed-Trim-UCB, to maintain robustness against the Byzantine clients (Section 4).
  • We derive $(\beta,\delta)$-PAC sample complexity results and a high probability $O(1)$ upper bound on the cumulative regret for the Fed-MoM-UCB algorithm (Section 5).
  • We conduct detailed experiments to compare Fed-MoM-UCB’s performance against Fed2-UCB in [24] and Fed-Trim-UCB under Byzantine attacks and to investigate the effects of different parameters on the performance (Section 6).
2 Related Work
Table 1: Comparison of different works on FMAB problems
@@ 2 Related Work
HetoFedBandit [30] $\times$ $\widetilde{O}(d\sqrt{MNT})$$\times$
Byzantine-UCB [31] $\times$ $\widetilde{O}(T^{3/4})$$\times$
@@
2.1 Federated Learning
In recent years, FL has experienced significant advancements, with particular focus on addressing key challenges such as privacy, communication costs, and data heterogeneity [34, 35]. The work by [36] introduces MiTFed, an innovative framework that leverages blockchain technology and smart aggregation to enhance robustness against Distributed Denial of Service (DDoS) attacks, ensuring privacy and facilitating collaboration in a decentralized context. Another study by [37] tackles the issues of communication overhead and data heterogeneity—two major obstacles in FL as identified by [38]. The authors propose the COTAF algorithm, which preserves the convergence properties of local Stochastic Gradient Descent (SGD) while accommodating heterogeneous data across users. Additionally, [39] addresses data heterogeneity by quantifying the skewness of client data in a privacy-preserving manner, subsequently selecting clients with less skewed data using a bandit approach. Scalability of FL systems in edge networks is studied in [40], where a fully decentralized FL system is shown to achieve significant gains in scalability and adaptivity.
@@ 2.1 Federated Learning
Some other related recent works include [51] which investigate Byzantine-robust distributed online learning, showing that even with robust aggregation, Byzantine-robust online gradient descent incurs linear adversarial regret. However, in i.i.d. environments, they achieve sublinear stochastic regret using momentum. Other studies such as [52, 53] also provide sublinear regret results under i.i.d. environments, where [52] utilizes a trimmed mean estimator to achieve robustness and [53] uses tools from high-dimensional robust Gaussian mean estimation. [54] considers a different approach and propose a robust federated learning framework that integrates blockchain technology to defend against poisoning attacks by decentralizing the model aggregation process. They combine the blockchain technology with a majority voting system to validate updates.
2.2 Federated Multi-armed Bandits
@@
FMAB problems share common challenges with FL, and several studies have sought to address various FL aspects within the FMAB framework. Table 1 provides a comparison of different FMAB studies, outlining the specific challenges each addresses. Notably, [24] was among the first to introduce the FMAB framework, with a focus on data heterogeneity and communication costs. They achieve a regret bound of $O(\log(T))$. Another significant work is by [31], which explores a federated linear bandit scenario involving Byzantine clients, albeit without considering data heterogeneity. Their main objective is to enhance the performance of each uncorrupted client by minimizing a metric they refer to as robust regret, using a robust geometric median aggregation. They manage to achieve a sublinear $\widetilde{O}(T^{3/4})$ regret, without data heterogeneity. [30] also explores the federated linear bandit setting, but focuses on data heterogeneity, rather than Byzantine robustness. Their algorithm introduces a clustering approach that enables collaboration among clients with similar reward models. They obtain a regret bound of $\tilde{O}(d\sqrt{MNT})$ where $d$ is the dimension of the action space, $N$ is the number of clusters and $M$ is the number of clients. Their approach offers a fully personalized solution, positioning it at the other end of the spectrum compared to [24]. A related work [32] addresses the FMAB problem with a focus on both data heterogeneity and Byzantine robustness. Unlike our approach and that of [30], their concept of data heterogeneity is based on varying arm availability across clients, rather than differences in arm means. This study specifically targets the best arm identification problem, achieving high-probability identification guarantees. Their algorithm combines a local successive elimination phase with a global best arm identification phase. To ensure Byzantine robustness, they employ a majority voting mechanism with trimming to filter out adversarial influences. Recently, global reward maximization by utilizing partial distributed feedback is studied by [55] within the context of linear bandits, while also aiming differential privacy. Another study by [33] investigates the federated linear bandits problem and presents an algorithm that employs layered successive screening. They separately address the scenarios of heteroscedastic noise and adversarial attacks, providing a regret bound that scales linearly with the corruption budget in the adversarial setting. They achieve a regret bound of $\widetilde{O}(\sqrt{dT}+dC_{p})$ where $d$ is the dimension of the action space and $C_{p}$ is the total corruption budget of the adversary. Each study mentioned adopts a distinct approach, combining various aspects. In our work, we address both $(\beta,\delta)$-PAC BAI and regret minimization amidst heterogeneous data and Byzantine attacks, while also factoring in communication costs.
Figure 1: Illustration of the motivating example.
2.3 Motivating Example
Consider the problem of selecting effective treatments across various clinics or regions, coordinated by a central medical research institution as shown in Fig. 1. Here, different treatment options can be represented as arms in a MAB problem. The problem is inherently federated, as data is collected from multiple clinics, with privacy regulations limiting the extent to which patient data can be shared. Additionally, model heterogeneity arises due to diverse patient demographics across clinics, leading to variations in treatment outcomes. Robustness against Byzantine clients is essential in this context, as erroneous data—whether from equipment malfunction or deliberate tampering—could result in the selection of a suboptimal or even harmful treatment. Instances of clinical malpractice in data handling are well-documented in the literature [56].
In times of crisis, such as during a novel pandemic, quickly identifying a satisfactory treatment option is crucial to improving public health, framing this as a $(\beta,\delta)$-PAC best arm identification problem. At the same time, minimizing regret is also important, as suboptimal treatments can have negative consequences for patient health. Motivated by these challenges, our study investigates the Federated Multi-Armed Bandit framework under Byzantine attacks, with a dual emphasis on best arm identification (BAI) and regret minimization.
3 Problem Formulation
3.1 Federated MAB Setup
@@
We consider a federated learning system with a server and a pool of clients that can be recruited by the server. There are $K$ arms indexed by the set $[K]\coloneqq\{1,2,...,K\}$. Arm $k$ has an unknown global mean reward $\mu_{k}$. We call $\boldsymbol{\mu}:=(\mu_{k})_{k\in[K]}$ the global model. We denote the global optimal arm by $k^{*}\coloneqq\operatorname*{argmax}_{k\in[K]}\mu_{k}$ and assume that it is unique. Clients cannot directly interact with the global model, but instead, they interact with their local models, which are stochastic realizations of the global model. In the local model of client $m$, the local mean reward of arm $k$ is generated as $\mu_{m,k}=\mu_{k}+\xi_{m,k}$, where $\xi_{m,k}$ is the $\sigma_{c}$-subgaussian model noise. For another illustrative example of the model noise, consider a recommender system in which the server’s objective is to pinpoint the item enjoying the highest global popularity, aiming to maximize profits. Due to the server’s lack of direct access to global popularities, it relies on feedback from selected clients. Nonetheless, the popularities observed by clients at a local level may diverge from the global trends. The random variable $\xi_{m,k}$ represents this discrepancy. We assume that $\sigma_{c}$ is known by the server.
@@
$X_{m,k}(t)=\mu_{m,k}+\zeta_{m,t}=\mu_{k}+\xi_{m,k}+\zeta_{m,t}~{}.$ (1)
In the above display, $\zeta_{m,t}$ is the sampling noise that is drawn from a fixed $\sigma$-subgaussian distribution independent of other clients and rounds. We assume that $\sigma$ is known by the server. Given $\beta>0$ and $\delta\in(0,1)$, the server and the clients’ main joint goal is to identify an arm whose global mean is within $\beta$ of the global optimal arm with at least $1-\delta$ probability. In the bandit literature, this is also known as $(\beta,\delta)$-PAC BAI [9]. Let $\Delta_{k}:=\mu_{k^{*}}-\mu_{k}$ denote the suboptimality gap for arm $k$. For $\beta\geq 0$, let $\mathcal{K}_{\beta}:=\{k\in[K]~{}|~{}\Delta_{k}\leq\beta\}$ denote the set of $\beta$-optimal arms. Let $\neg\mathcal{K}_{\beta}\coloneqq[K]\setminus\mathcal{K}_{\beta}$ denote the set of $\beta$-suboptimal arms. We define our objective as finding an arm $k\in\mathcal{K}_{\beta}$. Note that when $\beta<\min_{k\in[K]\setminus\{k^{*}\}}\Delta_{k}$, this coincides with finding the optimal arm.
While $\zeta_{m,t}$ is sampled at each round $t$, $\xi_{m,k}$ is sampled only once at the beginning, and $\mu_{m,k}$ is fixed from thereon. Equivalently, we can view $X_{m,k}(t)$ as a sample from an unknown distribution $\nu_{m,k}$. We call $\boldsymbol{\nu}_{m}:=(\nu_{m,k})_{k\in[K]}$ the local model of client $m$. For different clients $m$ and $n$, we have $\mu_{m,k}\neq\mu_{n,k}$ in general. Since the local mean rewards $\mu_{m,k}$ are different from the global mean rewards $\mu_{k}$, a client $m$ may not agree with the global optimal arm, i.e., $k^{*}\notin\operatorname*{argmax}_{k\in[K]}\mu_{m,k}$. In such a case, it is a hopeless task to identify the global optimal arm by relying on the local observations of a single client. However, the structure of the model noise suggest that by recruiting many clients, and averaging over their samples, one should be able to get an accurate estimate of the global mean rewards. This idea is exploited in [24], where the interaction between the server and the clients proceeds in phases indexed by $p\geq 1$. At the end of each phase, clients send updates to the server, denoted by $U_{m,k}(p)$, which are estimates of the local mean rewards of arms. The server, then aggregates these updates to discard suboptimal arms and informs the clients. After this, the clients refine their updates by sampling the remaining arms in the next phase. This interaction continues until $k^{*}$ is identified.
3.2 Byzantine Clients
@@
In our study, we focus on the challenge of identifying $\beta$-optimal arms, namely, those arms residing within a $\beta$-proximity of the global optimal arm. To address the robustness concern, we introduce a median-of-means (MoM) aggregator for the server, detailed further in Section 4. It will be impossible to learn $\beta$-optimal arms when all recruited clients are Byzantine. In order to obtain meaningful results, we make the following assumption on the ratio of the Byzantine clients. This assumption is standard in the robust statistics literature, where it is assumed that an upper bound on the fraction of corrupted samples is known [57, 58, 59, 32]. Similar assumptions are common in the federated learning literature addressing Byzantine attacks. For instance, [32, 51, 31, 53] assume an estimated upper bound on the number of Byzantine clients, while [33] assumes a bounded total corruption.
@@
3.3 Regret
In addition to $\beta$-optimal arm identification, we also want our algorithm to avoid sampling $\beta$-suboptimal arms unnecessarily, until all $\beta$-suboptimal arms are eliminated. We measure the performance of the interaction protocol between the server and the clients by introducing the notion of robust federated regret. Let $I_{t}^{m}$ denote the arm selected by client $m$ at round $t$. We denote by $\mathcal{M}_{t}^{\beta}\subseteq\mathcal{M}$ the set of clients who plays an arm $k\in\neg\mathcal{K}_{\beta}$ at round $t$. Robust federated regret defined below accounts for the cumulative loss of all clients with respect to $\beta$-optimal arms by considering the communication cost between the server and clients.